Ninjoy:

If x*x+2x-35=0 then x=?

alones:

Simplify both sides of the equation. \( x^2+2x−35=0\) Since \(xx=x^2\)

alones:

Than factor left side \((x−5)(x+7)=0\)

alones:

now set factors so they would equal \(0\) \(x−5=0~ or~ x+7=0\)

alones:

\(x=5 ~or~ x=−7\) would be your final answer (: